Find the values of the trigonometric functions of t from the given information.
cos t = − 12/13
Terminal point of t is in Quadrant III
sin t =
tan t =
csc t =
sec t =
cot t =
Please be descirptive on how this is solved. Thank you

Answer by lwsshak3(11628)

Find the values of the trigonometric functions of t from the given information.
cos t = − 12/13
Terminal point of t is in Quadrant III
Given information show that you are working with a 5-12-13 reference right triangle in quadrant III in which cos<0, sin<0
sin t =-5/13
tan t =5/12
csc t =-13/5
sec t =-13/12
cot t =12/5